Assignment 1: Problem 8 Previous Problem Problem List Next Problem (1 point) A statistical investigation is conducted to see if the mean of a population of values is equal to 168, or Ho 168 HA: 168 A random sample of n 57 values is taken from this population of values, from which the mean X and the standard deviation S were computed. From these, the test statistic was found to equal -0.23728. (a) Find the P-value. Use at least four decimals. P 1.1866 (b) Choose the correct interpretation of the P-value found in (a). OA. The probability computed in (a) is the probability of rejecting the null hypothesis, assuming that the null hypothesis is true. B. The probability computed in (a) is the chance of another sample producing stronger evidence against the null hypothesis, assuming that the null hypothesis is true OC. The probability in (a) is the probability of concluding the null hypothesis to be true. OD. The probability in (a) is the probability of rejecting the null hypothesis.
Answer #1

a) This is a two tailed test since alternative hypothesis contains "\neq" sign.

Degrees of freedom = n - 1 = 57 - 1 = 56

Hence,

P - value = 2*P(t56 < -0.23728) = 2*0.40665 = 0.8133

b) Option B is correct.